OpenStudy (anonymous):

can someone please help me find the inverse of f(x) = -3x + 21. and determine whether the inverse is a function, and state its domain and range.

OpenStudy (anonymous):

\[a)f^{-1} (x)=-\frac{ 1 }{ 3 }x+7\] ^not a function^ \[b)f^{-1} (x)=\frac{ -x+7 }{ 3 }\] ^not a function^ \[c)f^{-1} (x)=-\frac{ 1 }{ 3 }x+7\] ^^function; D: all real numbers; R: all real numbers.^^ \[d)f^{-1} (x)=\frac{ -x+7 }{ 3 }\] ^^function; D: all real numbers; R: all real numbers.^^